OpenStudy (anonymous):

. Simplify the following expression, and rewrite it in an equivalent form with positive exponents. -18xy^5/54x^3y^6

OpenStudy (anonymous):

you can simplify the numerical part (the "-18" and "54") like a normal fraction (-18/54)... look for common factors in numerator and denominator and cancel them.

OpenStudy (anonymous):

3 is common in both

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

right... actually... if you notice, 18 is common in both.... 3 x 18 = 54 So you could cancel out 18 from the top, leaving 1, and from the bottom, leaving 3 (oh, you beat me to it...)

OpenStudy (anonymous):

So, it's the same thing for the variables. You have x on top and x^3 on bottom

OpenStudy (anonymous):

cancel to leave x^2 on bottom.

OpenStudy (anonymous):

y^5 on top, and y^6 on bottom... cancel to leave just y on the bottom

OpenStudy (anonymous):

overall, looks like -1/(x^2y)

OpenStudy (anonymous):

would it be 1y or 11y

OpenStudy (anonymous):

I don't understand...\[\frac{ y ^{5} }{ y ^{6} } = \frac{ 1 }{ y }\]

OpenStudy (anonymous):

-1/3x^2y

OpenStudy (anonymous):

this is the answer

OpenStudy (anonymous):

\[-\frac{ 1 }{ 3x ^{2}y }\]

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

yes, I got confused when you asked something about 11y

OpenStudy (anonymous):

yea i get confused on if you add or subtract the variables

OpenStudy (anonymous):

oh, ok.... if it's a fraction with variables with exponents, then think of it like this for like y^5/y^6: \[\frac{ (y)(y)(y)(y)(y) }{ (y)(y)(y)(y)(y)(y) }\]

OpenStudy (anonymous):

so you just think to yourself... "I have 5 y's on top, and 6 on the bottom, so I will cross out 5 from each, leaving just one y on the bottom." You could also describe that as "subtracting the exponents", as in, you will subtract the exponent 5 on the top y's from the exponent 6 on the bottom y's, leaving just y^1 on the bottom, since 6-5 is 1. This is why the question told you to answer with POSITIVE exponents... You COULD (if you wanted) cancel the other way, subtracting the 6 exponent on the bottom from the 5 on the top, leaving -1 on the top.... and y^(-1) just happens to be = 1/y^1 so you get the same thing, but expressing the answer as POSITIVE exponents is "cleaner"

OpenStudy (anonymous):

so as positive it is no - answers

OpenStudy (anonymous):

if you had ended with x^(-2)y^-1, it would have been correct but wrong form

OpenStudy (anonymous):

correct value, but using negative exponents

OpenStudy (anonymous):

so the answers is - also this is were i really get lost.

OpenStudy (anonymous):

Oh!!! no... different!!

OpenStudy (anonymous):

It's fine that the whole thing is negative... they just didn't want you to answer like the following:

OpenStudy (anonymous):

\[-\frac{ x ^{-2}y ^{-1} }{ 3 }\]

OpenStudy (anonymous):

can you help me with another one

OpenStudy (anonymous):

Short Answer: Simplify the following expression, and rewrite it in an equivalent form with positive exponents. (5x-1)-3

OpenStudy (anonymous):

That is equivalent to what we had above, but using negative exponents... if you had done it that way, you "move" the x^(-2) to the bottom by multiplying top and bottom by x^2, cancelling the x^(-2) on top and leaving x^2 on bottom...

OpenStudy (anonymous):

did you leave out the exponent part of that 2nd problem?

OpenStudy (anonymous):

or is that written correctly?

OpenStudy (anonymous):

(5x^-1)^-3

OpenStudy (anonymous):

125/x^3 is this right

OpenStudy (anonymous):

So, that is really like: \[(5\frac{ 1 }{ x })^{-3} = \frac{ 1 }{ (\frac{ 5 }{ x })^{3} }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me double check that...

OpenStudy (anonymous):

easier: 5x^-1 is just 5/x so (5x^-1)^(-3) is (5/x)^-3 which is equal to (x/5)^3

OpenStudy (anonymous):

So it would be x^3 / 125 I think that's it, anyway... bad to hurry, need to make sure you don't flip the fraction over too many times while playing with that negative exponent

OpenStudy (anonymous):

ok i switch it around

OpenStudy (anonymous):

"raising to the negative 3rd power" means the same as "cubing and inverting" \[x ^{-3} = \frac{ 1 }{ x }\]

OpenStudy (anonymous):

ugg,,, that was supposed to be\[\frac{ 1 }{ x ^{3} }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I just said that as a way of saying that neg exponents are sometimes confusing, but if you start to get confused, just rewrite as a fraction and "move" the neg exponent variable across the fraction (either from top to bottom or bottom to top) to make it a positive exponent

OpenStudy (anonymous):

ok thanks for everything

OpenStudy (anonymous):

no problem :) These are annoying to type with all the exponents... :)